A pendulum is raised 0.7m above the floor, is released and then swings freely. Neglecting all frictional forces, what is the velocity of the pendulum when it is 0.25 m above the floor?
1
Expert's answer
2020-05-15T16:22:56-0400

According to law of conservation of energy, the potential energy between 0.7 and 0.25 m converted to the kinetic energy:


"mg(H-h)=\frac{1}{2}mv^2,\\
\space\\
v=\sqrt{2g(H-h)}=2.97\text{ m/s}."